Blossom Co. purchased equipment for $600,400 which was estimated to have a useful life of 10 years with a salvage value of $8,400 at the end of that time. Depreciation has been entered for 7 years on a straight-line basis. In 2018, it is determined that the total estimated life should be 15 years with a salvage value of $4,600 at the end of that time.

(a) Prepare the entry (if any) to correct the prior years’ depreciation.
(b) Prepare the entry to record depreciation for 2018.


(Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)
Answer #1
a
Account titles and explanation Debit Credit
No Entry 0
     No Entry 0
b
Account titles and explanation Debit Credit
Depreciation expense 22675
       Accumulated Depreciation-Equipment 22675
Workings:
Cost 600400
Less: Salvage value 8400
Depreciable cost 592000
Divide by Useful life 10
Annual depreciation 59200
Accumulated depreciation 414400 =59200*7
Book value 186000 =600400-414400
Less: Salvage value 4600
Remaining Depreciable cost 181400
Divide by remaining Useful life 8 =15-7
Annual depreciation 22675